# From Friendship Networks to Classroom Dynamics: Leveraging Neural Networks, Instrumental Variable and Genetic Algorithms for Optimal Educational Outcomes

### v4

@@ 3 Data description and data notation
We use the publicly available China Education Panel Survey (CEPS) data provided by the National Survey Research Center at Renmin University of China (NSRC). CEPS conducted surveys on 7,649 students, along with their family members and teachers, across 179 classes. For classes selected to participate, all students were surveyed. We use $N$ to denote the number of students in a single classroom. If it is for a particular classroom $c$ (from school from $s$), we use $N_{c}$ ($N_{cs}$) to denote the class size.
CEPS collected information from students and their family members regarding the students’ 6th-grade details, for example: class rank (in quantile), academic subject interest, extracurricular hobbies, etc. This aspect of the survey is particularly useful for our project. When assigning students to classrooms upon school entry, the principal only has access to their 6th-grade (primary school) information. Therefore, we exclusively use students’ 6th-grade data and other predetermined characteristics (e.g. gender and locality), collectively denoted as $X$, to predict friendship formation. One particular variable in $X$ that we repeatedly utilize in different steps of our methodology is students’ 6th-grade class quantile, denoted as $z$. If we need to emphasize the 6th-grade quantile of all students from classroom $c$ (or a particular student $i$ from classroom $c$ and school $s$), we use $z_{c}$ (or $z_{ics}$) to denote that information.
@@
Students were asked to provide aggregated relational data (ARD) of their friends. The linkage data, though surveyed in section C20 as depicted in Figure 6 in Appendix B, is redacted from CEPS data. Instead, ARD of the friendship network (See section C21 from Figure 6) is provided. Therefore, we have information on ‘out of the five best friends of a student, how many of them                           ?’ where                            can be filled by any question from section C21. We use ARD, denoted as $A_{f}$, as the response variable for PeerNN to uncover friendship formation patterns. $A_{f}$ has size $(N\times 10)$ and records each student’s responses to 10 distinct questions about their friends’ gender, locality, relationship status, etc. Each question only has three possible answers: ‘none’, ‘one or two’, or ‘most of them’. For example, one student may report none of his friends are female, most of his friends are local, and one or two of his friends are in a relationship, etc. Apart from $A_{f}$, each student’s own information regarding these 10 questions is also available in CEPS. The answer to each question is binary (yes or no). For example, one student is female, local, and in a relationship, etc. We denote students’ own information as $A_{s}$, which has the same dimension as $A_{f}$. We use $A_{fiq}$ to denote ARD information provided by student $i$ about question $q$ and $A_{sq}$ to denote all students’ own characteristics for question $q$.
@@
Upon entering their respective secondary schools, 5860 (from 139 classrooms) out of 7649 students (from 179 classrooms) were randomly assigned to classrooms. This natural experiment setup helps generate exogenous variation in data. Exploiting this data-generating property, we construct instrumental variables for causal inference analysis. To keep our input data consistent for all three steps, we use these 5860 students’ data (1) as the training set for PeerNN, (2) to estimate peer effect, and (3) to design classroom assignments. The rest of the 1789 students who are non-randomly assigned to classrooms are used as test data for PeerNN.
@@ 4 Model setup
Problem formulation: The ultimate goal of this work is to devise a class assignment policy that maximizes peer effect. The school principal’s objective function is defined as the following
$\max_{\pi\in\Pi}\frac{1}{\sum_{c\in C}N_{c}}\sum_{c\in C}\beta\mathbf{1_{N_{c}}^{\top}}\Omega_{c}(\pi)z_{c}(\pi)$ (1)
where $\pi$ is a policy from a set of policies, $\Pi$, that satisfy mandatory institutional constraints such as balanced gender ratio across classrooms, $C$ is a set of consecutive integers $\{1,2,\dots,|C|\}$ indicating the indices of classrooms in a school and size of $C$ (i.e. $|C|$) is a known fixed number, $\mathbf{1_{N_{c}}}$ denotes a vector of ones of length $N_{c}$, $\{N_{c}\}_{c\in C}$ is a set of integers numbers indicating the prespecified classroom sizes for each classroom $c$, $\Omega_{c}$ is the adjacency matrix of classroom $c$ and $z_{c}$ is the 6th-grade class quantile of all students in the classroom $c$. Both $\Omega_{c}$ and $z_{c}$ are functions of $\pi$ since $\pi$ determines which students are assigned to classroom $c$. $\beta$ is the peer effect parameter: if a student’s friends’ average 6th-grade class quantile improves by 10%, then the student’s cognitive ability test score improves by $\frac{\beta}{10}$ point, holding everything else fixed. We are particularly interested in the sign of $\beta$.
As shown by Figure 1, we solve the class assignment optimization problem with the following three steps: (i) fitting/predicting friendship formation probability, $\Omega_{c}$, using $X$, (ii) measuring peer effect, $\beta$, with an instrumental variable approach, and (iii) heuristically searching a class assignment policy, $\pi$, that maximizes expression (1) with genetic algorithm.
Figure 1: The three-step framework that takes in predetermined characteristics of students and outputs an optimal class assignment policy.
@@ 4.1 Friendship formation neural network
We build an interpretable neural network architecture, termed PeerNN, to predict friendship formation with ARD. PeerNN can be divided into four stages: (1) friendship market, (2) linearized propensity score, (3) normalization, and (4) loss function. Next, we describe the four stages along with their interpretation. For notation convenience, the description pertains to a single classroom setting with class size $N$.
@@
We encode our raw data $X$ to a lower dimensional embedding $\sigma(X)$ with a fully connected hidden layer that contains ten nodes, $\sigma(X)$ is an $(N\times 10)$ matrix. This allows students to form friendship preferences based on latent representations of $X$ (i.e. $\sigma(X)$). Stage 0 serves as a feature extraction step for Stages 1-4 and is estimated as part of PeerNN.
@@
Students’ preference parameters $\delta$ (dimension $(N\times 10)$) are functions of $\sigma(X)$. The ith row of $\delta$ matrix, $\delta_{i}$ is the preference parameters of student $i$. $\delta_{ik}$ indicates how much student $i$ prefers to have a friend with a higher value of latent characteristic $k$. Example C.2 illustrates the concept of preference parameters. We use two fully connected layers to model preference parameters $\delta(X)$:
@@
Stage 2: Linearized propensity score
We obtain the outer product of $\delta(X)$ and $\sigma(X)$ and denote the outer product as $\Upsilon$, an $N\times N$ matrix, where $\Upsilon_{ij}=\sum_{k=1}^{K}\delta_{ik}\sigma_{jk}=(\delta\cdot\sigma^{\prime})_{ij}$. $\Upsilon_{ij}$ corresponds to a concept in the structural applied microeconomics literature (McFadden, 1974; Train, 2009): the deterministic part of the utility function for student $i$ picking student $j$ as best friend. Alternatively, we can interpret it as the linearized propensity score in causal inference.
@@
We manually set $\Upsilon_{ii}$ as negative infinity since students cannot make friends with themselves. Utility for student $i$ to consider student $j$ as a best friend is subject to a utility shock $\xi_{ij}\overset{iid}{\sim}$ Gumbel distribution: $U_{ij}=\Upsilon_{ij}+\xi_{ij}$. Student $i$ picks his best friend based on who gives him the highest utility, $\operatorname*{\arg\!\max}_{j}U_{ij}$. Knowing the joint distribution of $\{\xi_{ij^{\prime}}\}_{j^{\prime}=1}^{N}$, the probability of student $i$ picking student $j$ as best friend is given by softmax function $P(\operatorname*{\arg\!\max}_{j^{{}^{\prime}}}U_{ij^{{}^{\prime}}}=j)=\frac{\exp{(\Upsilon_{ij})}}{\sum_{j^{\prime}}^{N}\exp{(\Upsilon_{ij^{\prime}})}}:=\Omega_{ij}$. The adjacency-probability matrix $\Omega$ is obtained by applying the softmax function on $\Upsilon$ row-wise. Example C.3 in Appendix C demonstrates how Stage 2 and 3 work.
@@
Two issues arise. Due to CEPS’s survey design, an exact mapping from $(\Omega_{i},A_{sq},B_{i})$ to $A_{fiq}$ requires nonlinear operations, rendering MSE to have no closed form. Moreover, sampling friends without replacement based on the probability vector $\Omega_{i}$ is intractable in backpropagation. See appendix D.1 for more details on these issues. To address the two problems, we make two compromises: (1) approximating nonlinear operations with multiple linear functions, and (2) drawing friends with replacement. We can decompose MSE into ${\rm Bias}^{2}$ and ${\rm Var}$. With the two compromises, we derive closed-form expressions for ${\rm Bias}^{2}$ and ${\rm Var}$, exponentially reducing the computational complexities and making the backpropagation process tractable. The description of the two compromises and the derivations of the closed-form expressions are shown in Appendix D.1.
@@
Additional notes on loss function: When running the empirical analysis, we find it helpful to downweigh variance. The benefit of downweighing variance can be intuitively rationalized by considering network density. We defer the explanation to Appendix D.2 as it does not alter the main idea of why the loss function makes sense. The modified loss function can be written as ${{\rm Bias}}^{2}+\mu{{\rm Var}}+\kappa H(X)+\lambda T(X)$, where $(\mu,\kappa,\lambda)$ are hyperparameters. We can avoid cross-validation for tuning hyperparameters by examining whether $\Omega$ satisfies commonly known properties of social networks like gender homophily, presence of central nodes, and clustering. Such examination of $\Omega$’s interpretability can help pin down a range of reasonable values of $(\mu,\kappa,\lambda)$. We substantiate our claim in Appendix D.3.

### v5

@@ 2 Notation and data description
We use the publicly available China Education Panel Survey (CEPS) data provided by the National Survey Research Center at Renmin University of China (NSRC). CEPS conducted surveys on 7,649 students, along with their family members and teachers, across 179 classes. For classes that are selected to participate, all students were surveyed.
2.1 Classroom size and school size
We use $N$ to denote the number of students in a single classroom. If we want to emphasize classroom $c$ (or a classroom $c$ from school from $s$), we use $N_{c}$ (or $N_{cs}$) to denote the class size. Additionally, we use $N_{s}$ to denote the number of students in school $s$.
2.2 Predetermined characteristics
CEPS collected information from students and their family members regarding the students’ 6th-grade details, for example: class rank (in quantile), academic subject interest, extracurricular hobbies, etc.
When assigning students to classrooms upon school entry, the principal only has access to their information predetermined before the students enter middle school. Therefore, we exclusively use these predetermined variables (e.g. gender and locality), collectively denoted as $X$, to predict friendship formation. Occasionally, we use the notation $X_{c}$ (or $X_{s}$) for the predetermined characteristics of all students from a classroom $c$ (or from a school $s$).
One particular variable in $X$ that we repeatedly use in different steps of our methodology is students’ 6th-grade class quantile, denoted as $z$. If we need to emphasize the 6th-grade quantile of all students from classroom $c$ (or of a particular student $i$ from classroom $c$ and school $s$ or of all students from school $s$), we use $z_{c}$ (or $z_{ics}$ or $z_{s}$) to denote that information.
2.3 Aggregated relational data
@@
Students were asked to provide aggregated relational data (ARD) of their friends. The linkage data, though surveyed in section C20 as depicted in Figure 10 in Appendix A, is redacted from CEPS data. Instead, ARD of the friendship network (See section C21 from Figure 10) is provided. Therefore, we have information on ‘out of the five best friends of a student, how many of them                           ?’ where                            can be filled by any question from section C21.
We use ARD, denoted as $A_{f}$, as the response variable for PeerNN to uncover friendship formation patterns. $A_{f}$ has size $(N\times 10)$ and records each student’s responses to 10 distinct questions about their friends’ gender, locality, relationship status, etc. Each question only has three possible answers: ‘none’, ‘one or two’, or ‘most of them’. For example, one student may report none of his friends are female, most of his friends are local, and one or two of his friends are in a relationship, etc.
Apart from $A_{f}$, each student’s own information regarding these 10 questions is also available in CEPS. The answer to each question is binary (yes or no). For example, one student is female, local, and in a relationship, etc. We denote students’ own information as $A_{s}$, which has the same dimension as $A_{f}$. We use $A_{fiq}$ to denote ARD information provided by student $i$ about question $q$ and $A_{sq}$ to denote all students’ own characteristics for question $q$.
2.4 Train and test sets
@@
Upon entering their respective secondary schools, 5860 (from 139 classrooms) out of 7649 students (from 179 classrooms) were randomly assigned to classrooms, as indicated by a survey question in CEPS. This natural experiment setup helps generate exogenous variation in data. Exploiting this data-generating property, we construct instrumental variables for causal inference analysis. To keep our input data consistent for all three steps, we use these 5860 students’ data (1) as the training set for PeerNN, (2) to estimate peer effect, and (3) to design classroom assignments. The rest of the 1789 students who are non-randomly assigned to classrooms are used as test data for PeerNN.
2.5 Data summary
Table 1 presents students’ demographic and background characteristics in the full, training, and test datasets. In the full sample, the average class size is approximately 45 students, and the average student age is 167.57 months (around 14 years). 47.2% of the students are female, 50.4% hold rural Hukou, and 8.8% belong to non-Han ethnic groups. The average 6th-grade academic quantile is 66%. The average father’s and mother’s education level is 4, corresponding to a technical secondary school or technical school degree—above junior high school but below senior high school.
In the training dataset, the average class size is 46 students, with the average age of 167.57 months (14 years). Among these students, 48% are female, 48.5% have rural Hukou, and 8.7% are non-Han. Parental education levels remain consistent with those in the full sample. In the test dataset, the average age remains the same, but the gender and background composition differs slightly compared to the training. Female students make up a lower proportion (44.7%), while the proportion with rural Hukou is higher at 56.5%, and 9.3% are non-Han. The test data also show a higher average class size (47 students) and slightly lower average education levels for both parents compared to the training dataset.
Full Train Test
Mean S.D. Mean S.D. Mean S.D.
Age 167.568 7.939 167.565 7.790 167.575 8.410
Gender 0.472 0.499 0.480 0.500 0.447 0.497
Hukou 0.504 0.500 0.485 0.500 0.565 0.496
6th-grade quantile 0.660 0.233 0.660 0.236 0.662 0.223
Class size 45.057 11.529 45.682 11.869 47.285 10.243
Father’s education 4.329 2.024 4.419 2.061 4.035 1.871
Mother’s education 4.067 2.007 4.169 2.038 3.733 1.867
Ethnic nationality 0.088 0.283 0.087 0.281 0.093 0.290
Number of observations 7,649 5,860 1,789
Table 1: Summary statistics of the full dataset, train set and test set
2.6 Test for random assignment
The training sample is selected based on a survey question that asks the teachers whether the classes that they teach are randomly assigned. On top of that, we further conduct likelihood ratio tests to check whether class assignment is based on students’ academic rank and other characteristics. The intuition behind this approach is that if students are randomly assigned to classes within each school, then their individual characteristics should be independent of class assignment. In other words, the proportion of students with specific characteristics in each class should not differ systematically from the corresponding proportions at the school level.
To test this, we conduct a likelihood ratio test (LRT) based on the distribution of students’ specific characteristics. Let $N_{c}$ and $N_{s}$ denote the number of students in the class and school respectively and $n_{c}$ and $n_{s}$ denote the number of students with a specific characteristic in the class and school respectively. When we randomly draw a class in each school, the likelihood ratio is
$\text{LR}_{c}=\frac{\mathcal{L}_{0}}{\mathcal{L}_{1}}=\frac{\left(\frac{N_{c}}{N_{s}}\right)^{n_{c}}\left(1-\frac{N_{c}}{N_{s}}\right)^{n_{s}-n_{c}}}{\left(\frac{n_{c}}{n_{s}}\right)^{n_{c}}\left(1-\frac{n_{c}}{n_{s}}\right)^{n_{s}-n_{c}}},$
where $\mathcal{L}_{0}$ is the likelihood under the null hypothesis, where is proportion in each class should equal to the school-level proportion, and $\mathcal{L}_{1}$ is the likelihood under the alternative hypothesis.
Then, we use the likelihood ratio to compute the $\chi^{2}$ statistics by using $\chi^{2}_{c}=-2\log(\text{LR}_{c})$ and get the $\chi^{2}$ distribution. We perform this test using several student characteristics: 6th-grade class academic quantile, Hukou type (rural or not), Hukou location (local or not), gender, father’s education, and mother’s education. Figure 2 displays the test results, with the black vertical line indicating the critical value. Most of the $\chi^{2}_{c}$ statistics fall below the critical threshold, suggesting that the within-class proportions of these characteristics do not significantly differ from the school-level proportions.
Figure 2: Likelihood ratio tests
Furthermore, we conduct Pearson’s $\chi^{2}$ tests within schools for each variable, following Ammermueller and Pischke (2009) and Wang and Zhu (2021). The results are presented in Figure 20 in the Appendix. We also conduct more conventional balance tests in Appendix I. The results support that the classrooms in the training set are randomly assigned.
@@ 3 Model setup
There are $N_{s}$ incoming students from middle school $s$ indexed by $i\in\mathcal{I}:=\{1,2,\dots,N_{s}\}$. The principal needs to assign the $N_{s}$ students to $|C|$ classrooms indexed by $\{1,2,\dots,|C|\}$. We denote the set of indices for classrooms as $C:=\{1,2,\dots,|C|\}$. A classroom assignment policy, $\pi$, is a partition of $\mathcal{I}$ with partition cardinality $|C|$ like in Goulas and Monachou (2024). Each subset in $\pi$ (the indices of students assigned to one classroom) is denoted as $\pi_{c}$ where $c\in C$, meaning that $\pi=\{\pi_{c}\}_{c\in C}$. The principal knows how many students should be assigned to each classroom, hence, the cardinality of $\pi_{c}$ is fixed before the students are assigned to classroom $c$ and denoted as $N_{c}$. The ultimate goal of the principal is to devise a policy $\pi$ that maximizes peer effect. The objective function of the principal of middle school $s$ is
$\max_{\pi\in\Pi}\frac{1}{N_{s}}\sum_{c\in C}\mathbf{1_{N_{c}}^{\top}}\Omega_{c}(\pi,X_{s})z_{c}(\pi,z_{s})\beta\quad\text{where}\quad\Omega_{c}(\pi,X_{s})=f(X_{c}(\pi,X_{s}))$ (3.1)
where $\Pi$ is the set of all policies satisfying the cardinality constraints on both the number of classrooms $|C|$ and the size of each classroom $\{N_{c}\}_{c\in C}$. $\mathbf{1_{N_{c}}}$ denotes a vector of ones of length $N_{c}$.
$X_{c}$ (an $N_{c}\times p$ matrix) and $z_{c}$ (an $N_{c}$ column vector) are predetermined characteristics and the 6th-grade class quantile of all students in classroom $c$. Our model allows $z_{c}$ to be one of the columns of $X_{c}$ (or not). We formally define $X_{c}:=\{X_{s,i}\}_{i\in\pi_{c}}$ and $z_{c}:=\{z_{s,i}\}_{i\in\pi_{c}}$, where $X_{s}$ and $z_{s}$ do not depend on $\pi$ and have $N_{s}$ rows, $\pi_{c}$ depends on $\pi$ and determines which rows of $X_{s}$ and $z_{s}$ should be included in $X_{c}$ and $z_{c}$, respectively.
Step 1 overview $f:\mathbbm{R}^{N_{c}\times p}\to[0,1]^{N_{c}\times N_{c}}$ maps predetermined student characteristics $X_{c}$ to a $N_{c}\times N_{c}$ friendship intensity matrix of classroom $c$, $\Omega_{c}$. $f$ is interpreted as the friendship formation process. In Section 3.1, we build a neural network architecture, PeerNN, to learn $f$. PeerNN takes $X_{c}$ as input and predicts $\Omega_{c}$ (i.e., $f(X_{c})$) as an intermediate output. The loss function of PeerNN is based on the match between $\Omega_{c}$ and ARD provided by the students from classroom $c$.
Step 2 overview $\beta$ is the peer effect parameter in the regression equation we specify later in Section 3.2 with the following interpretation: if a student’s friends’ average 6th-grade class quantile improves by 10%, then the student’s cognitive ability test score improves by $\frac{\beta}{10}$ point, holding everything else fixed. We will relax the homogeneous $\beta$ assumption in Section 4.2.3.
Step 1-3 overview Solving (3.1) requires nonconventional approaches for two reasons: first, $\Pi$ is discrete, making the optimization combinatorial and NP-hard; second, $\Omega_{c}$ is a function of both $\pi$ and $X_{s}$, making it impossible to write it as function of only $\pi$ or recast (3) as a one-step feasible programming problem like in Bhattacharya (2009); Carrell et al. (2013). As shown by Figure 3, we solve (3.1) with the following three steps: (i) fitting/predicting network $\Omega_{c}$ with PeerNN (ii) estimating peer effect parameter $\beta$ with an instrumental variable, and (iii) heuristically searching for the optimal $\pi$, that maximizes (3.1) with GA.
Figure 3: The three-step framework that takes in predetermined characteristics of students and outputs the optimal $\pi$. All subscripts are dropped because we illustrate the framework with respect to one single classroom. The first step (PeerNN) takes in $X$ and predicts $\Omega$ as an intermediate output; the second step incorporates $\Omega$ into the regression equation and estimates $\beta$; the sign of $\beta$ determines the third step (GA)’s fitness function, GA iteratively plugs in PeerNN to search for the optimal $\pi$.
@@ 3.1 Friendship formation neural network
We build an interpretable neural network architecture, termed PeerNN, to predict friendship formation with ARD. PeerNN can be divided into four stages: (1) friendship market, (2) utility function, (3) normalization, and (4) loss function. We provide concrete examples for each stage in Appendix B. The description pertains to a single classroom. Since the classroom is fixed, we drop the subscript $c$ in Section 3.1 for a cleaner notation.
@@
We encode our raw data $X$ to a lower dimensional embedding $\sigma(X)$ with a fully connected hidden layer that contains ten nodes, $\sigma(X)$ is an $(N\times 10)$ matrix. Stage 0 serves as a feature extraction step for Stages 1-4 and is estimated as part of PeerNN.
@@
In friendship market, students supply themselves as their latent features, $\sigma(X)$. Students’ preference parameters $\delta$ (dimension $(N\times 10)$) are functions of $\sigma(X)$. The ith row of $\delta$ matrix, $\delta_{i}$ is the preference parameters of student $i$. $\delta_{ik}$ indicates how much student $i$ prefers to have a friend with a higher value of latent characteristic $k$. We use two fully connected layers to model preference parameters $\delta(X)$:
@@
Example B.1 illustrates the concept of preference parameters. We strongly encourage readers who are not familiar with machine learning/neural networks literature to refer to Appendix B for the intuition behind PeerNN.
Stage 2: Deterministic part of utility function
We obtain the outer product of $\delta(X)$ and $\sigma(X)$ and denote the outer product as $\Upsilon$, an $N\times N$ matrix, where $\Upsilon_{ij}=\sum_{k=1}^{K}\delta_{ik}\sigma_{jk}=(\delta\cdot\sigma^{\prime})_{ij}$. $\Upsilon_{ij}$ corresponds to a concept in the structural applied microeconomics literature (McFadden, 1974; Train, 2009): the deterministic part of the utility function for student $i$ picking student $j$ as best friend.
@@
We manually set $\Upsilon_{ii}$ as negative infinity since students cannot make friends with themselves. Utility for student $i$ to consider student $j$ as a best friend is subject to a utility shock $\xi_{ij}\overset{iid}{\sim}$ Gumbel distribution: $U_{ij}=\Upsilon_{ij}+\xi_{ij}$. Student $i$ picks his best friend based on who gives him the highest utility, $\operatorname*{arg\,max}_{j}U_{ij}$. Knowing the joint distribution of $\{\xi_{ij^{\prime}}\}_{j^{\prime}=1}^{N}$, the probability of student $i$ picking student $j$ as best friend is given by softmax function
$P(\operatorname*{arg\,max}_{j^{{}^{\prime}}}U_{ij^{{}^{\prime}}}=j)=\frac{\exp{(\Upsilon_{ij})}}{\sum_{j^{\prime}}^{N}\exp{(\Upsilon_{ij^{\prime}})}}:=\Omega_{ij}.$
$\Omega$ is obtained by applying the softmax function on $\Upsilon$ row-wise. This formulation interprets $\Omega_{i}$, ith-row of the friendship intensity matrix, as the probability student $i$ considers each classmate as a best friend. Example B.2 in Appendix B demonstrates how Stage 2 and 3 work.
@@
Two issues arise. Due to CEPS’s survey design explained in Section 2.3, an exact mapping from $(\Omega_{i},A_{sq},B_{i})$ to $A_{fiq}$ requires nonlinear operations, rendering MSE to have no closed form. Moreover, sampling friends without replacement based on the probability vector $\Omega_{i}$ is intractable in backpropagation. See appendix C.1 for more details on these issues. To address the two problems, we make two compromises: (1) approximating nonlinear operations with multiple linear functions, and (2) drawing friends with replacement. We can decompose MSE into ${\rm Bias}^{2}$ and ${\rm Var}$. With the two compromises, we derive closed-form expressions for ${\rm Bias}^{2}$ and ${\rm Var}$, exponentially reducing the computational complexities and making the backpropagation process tractable. The description of the two compromises and the derivations of the closed-form expressions are shown in Appendix C.1. Here we only present the closed-form expressions:
where $(j,k)$ entry of $\rm{Var}_{\Omega_{i}}=B_{i}\Omega_{ij}(1-\Omega_{ij})$ when $j=k$ and $(j,k)$ entry of $\rm{Var}_{\Omega_{i}}=-B_{i}\Omega_{ij}\Omega_{ik}$ when $j\neq k$. $Q$ denotes the number of ARD questions and $(a_{B_{i}},b_{B_{i}})$ are as follows. The explanation for these numerical values of $(a_{B_{i}},b_{B_{i}})$ is in Appendix C. Here, we tabulate the values of $(a_{B_{i}},b_{B_{i}})$.
$B_{i}$ 1 2 3 4 5
$b_{B_{i}}$ 1. 1.090 1.154 1.2 1.333
$a_{B_{i}}$ 1.5 0.727 0.654 0.5 0.4
@@
Additional notes on loss function: When running the empirical analysis, we find it helpful to downweigh variance. The benefit of downweighing variance can be intuitively rationalized by considering network density. We defer the explanation to Appendix C.2 as it does not alter the main idea of why the loss function makes sense. The modified loss function can be written as ${{\rm Bias}}^{2}+\mu{{\rm Var}}+\kappa H(X)+\lambda T(X)$, where $(\mu,\kappa,\lambda)$ are hyperparameters. We can avoid cross-validation for tuning hyperparameters by examining whether $\Omega$ satisfies commonly known properties of social networks like gender homophily, presence of central nodes, and clustering. Such examination of $\Omega$’s interpretability can help pin down a range of reasonable values of $(\mu,\kappa,\lambda)$. We substantiate our claim in Appendix C.3.